# VectorLiteRAG: Latency-Aware and Fine-Grained Resource Partitioning for Efficient RAG

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems combine vector similarity search with large language models (LLMs) to deliver accurate, context-aware responses. However, co-locating the vector retriever and the LLM on shared GPU infrastructure introduces significant challenges: vector search is memory- and I/O-intensive, while LLM inference demands high throughput and low latency. Naive resource sharing often leads to severe performance degradation, particularly under high request load or large index sizes.
We present VectorLiteRAG, a deployment-friendly RAG system that achieves latency-compliant inference without requiring additional hardware resources. VectorLiteRAG introduces a fine-grained GPU resource allocation mechanism based on detailed performance modeling and access pattern analysis. By estimating search latency and query hit rate distributions, it identifies an optimal index partitioning point across CPU and GPU tiers to minimize contention and maximize throughput.
Our evaluations show that VectorLiteRAG consistently expands the SLO-compliant request rate range across all tested configurations, including both small and large LLMs, and small and large vector databases compared to naive baselines and state-of-the-art alternatives. In the best case, VectorLiteRAGimproves the attainable SLO throughput by up to 1.5× without compromising generation quality or requiring additional compute resources.
@@ I Introduction
Retrieval-Augmented Generation has emerged as a powerful system in natural language processing, particularly for domain-specific question answering and information retrieval tasks [21, 9, 31]. Its key strength lies in combining parametric memory, encoded in the weights of a large language model, with non-parametric memory retrieved from an external knowledge corpus. Although parametric memory provides strong generalization, it is expensive to train and difficult to update. To mitigate this, RAG pipelines first perform similarity search using approximate nearest neighbor search (ANNS) algorithms to retrieve relevant documents from a large database. The retrieved documents are then fed into the LLM’s context to generate up-to-date and reliable responses.
@@
RAG frameworks [19, 23, 36] typically adopt heterogeneous hardware configurations, where vector retrieval is executed on CPUs and LLM generation is served by GPUs. This division is driven by workload characteristics: LLM inference requires massive matrix multiplications and benefits significantly from GPU acceleration, while retrieval has traditionally been seen as a lighter task suited for CPUs. Offloading retrieval to CPUs allows GPUs to be dedicated to the more compute-intensive generation phase. CPU-based vector search may be sufficient for small vector databases. However, as the dimensionality of the embeddings and the size of the dataset grow, retrieval becomes increasingly compute- and memory-bound. CPUs, with limited parallelism, narrower vector units, and lower memory bandwidth, struggle to handle high-throughput similarity search at scale.
@@
Although the retrieval operation is computationally lighter than the generation phase, it can still benefit significantly from GPU acceleration for two reasons: (1) GPUs feature wide and powerful vector units that enable highly parallelized distance computations, offering superior performance for similarity calculations on long embedding vectors. (2) The retrieval process involves scanning intermediate distance tensors to identify the closest data points in the vector space. These operations are typically implemented as memory lookups a task where GPUs outperform CPUs due to their vectorized memory access and higher I/O throughput.
@@
These compute and memory pressures create a resource tension between the retrieval and generation stages, especially as the vector database grows and CPU-based search fails to meet strict latency requirements. GPU memory is already constrained, with most of it reserved for model weights and KV cache for the LLM. Naively sharding the vector index across all GPUs can lead to memory contention and reduced overall throughput. Alternatively, assigning a disaggregated GPU for retrieval can prevent direct interference between stages, but may degrade overall system throughput by reducing the number of available LLM instances, in particular when models require multiple GPUs, enforcing rigid allocation schemes. Motivated by these challenges, this work explores a holistic approach to optimizing distributed RAG pipelines through joint resource allocation between vector search and LLM generation. We present VectorLiteRAG, a system that partitions the vector index between GPU and CPU based on query access patterns and LLM deployment configurations, aiming to maximize throughput while meeting latency targets by exploiting the compute power of GPUs across both stages of the RAG pipeline. By analytically modeling similarity search latency, we determine the smallest index portion that needs to be placed on the GPU to satisfy the latency requirement under a given system configuration. Accordingly, VectorLiteRAG  offers a latency-aware, throughput-optimized solution that requires no additional hardware resources. This approach is grounded in two key insights:
Access-Skew-Aware Data Layout. VectorLiteRAG  leverages a key characteristic of IVF-based retrieval systems, that query accesses exhibit skew across clusters. To take advantage of this, VectorLiteRAG  introduces an analytical model that determines the optimal partitioning point and corresponding layout for a multi-GPU system. While the coarse quantizer and cold clusters remain on the CPU, a small subset of hot clusters are cached and distributed across GPUs. The system allocates just enough hot clusters to the GPUs, avoiding both oversubscription of GPU resources during retrieval.
Inter/Intra-Query Variance-Aware Routing. Hit rates vary across queries and, within a single query, across devices. Static retrieval configurations, such as fixed search budgets per device, fail to capture this variance and often result in unnecessary allocation of GPU thread resources. VectorLiteRAG  addresses this by incorporating query and device-level awareness into its task routing and resource allocation. After identifying the most relevant clusters, retrieval tasks are dispatched to GPUs or CPUs based on their expected contribution to overall search task. Moreover, by monitoring query progress during retrieval, VectorLiteRAG  opportunistically forwards early-completing queries to the LLM stage, reducing synchronization delays caused by straggler queries and improving latency under batching.
@@
  • Identifies and models access skew in IVF-based retrieval systems and proposes a hit rate estimation method based on observed cluster access patterns.
  • Develops an analytical latency model that predicts retrieval time while accounting for inter-query hit rate variance. Using this model, VectorLiteRAGdetermines an optimal index partitioning point that satisfies latency targets.
  • Designs a distributed retrieval pipeline that adaptively allocates search tasks across CPUs and GPUs by exploiting inter-device hit rate variance, thereby improving efficiency and avoiding unnecessary GPU resource usage.
@@ II Retrieval Augmented Generation
The Retrieval-Augmented Generation pipeline combines information retrieval with LLM inference to enable efficient, context-sensitive response generation [21, 8, 7]. User queries are first transformed into vector embeddings using embedding models [32, 28, 39, 34]. These embeddings capture the semantics of the input and enable similarity search by comparing query vectors to a vector database constructed from the knowledge corpus.
Since exhaustive pairwise search is computationally infeasible at scale, large-scale vector retrieval relies on approximate nearest neighbor search to efficiently identify relevant documents. The retrieved vectors are mapped back to their corresponding documents, which are provided as additional context to the LLM alongside the original query.
@@ II-A Inverted List Index IVF
The Inverted File (IVF) index [43], by contrast, organizes the index as a hierarchical clustering structure. A subset of vectors is first clustered via k-means to obtain centroids. Then, each database vector is assigned to the closest centroid, forming an inverted list. This structure narrows the search space using only centroid metadata, resulting in low memory overhead and high scalability. As such, IVF is widely adopted and studied in retrieval systems for large knowledge corpora [10, 33, 40, 5, 20, 13].
To further reduce memory usage, quantization techniques are applied on top of IVF. Scalar quantization (SQ) reduces each vector element to a smaller numerical type (e.g., float32 to int8), offering simplicity but limited compression. For higher compression ratios, product quantization (PQ) [12] is commonly used.
@@ II-B Search Operation in IVF Index
Next, a distance lookup table is constructed. Since each vector is quantized into discrete sub-vector codes, each code maps to a representative value, trained and stored in the codebook. By pre-computing distances between the query vector and these representative values, the system avoids computing full distances to every vector. During the scan stage, these LUTs are used to accumulate approximate distances and retrieve the top-k nearest vectors.
@@
Motivated by their superior latency-performance trade-off, we adopt fast scanning in our system to enable efficient and low-latency vector retrieval. However, despite the SIMD capabilities of modern CPUs, CPU-based search can still become a bottleneck, ultimately degrading the responsiveness of the RAG system.
@@ III Challenges and Opportunities in RAG serving
While fast scan indexes significantly improve the latency of vector similarity search on CPUs, GPU-based retrieval can offer even greater speedups, thanks to their wider vector processing units and higher memory bandwidth. As shown in Figure 4 (left), GPU-accelerated IVF search can outperform fast scan methods by nearly an order of magnitude.
Thus, offloading retrieval to the GPU becomes appealing for large-scale vector databases where CPU-based search remains a bottleneck. However, this comes with a fundamental trade-off: GPU memory is already heavily utilized by LLMs, particularly for storing KV cache and model weights. Allocating additional memory for the vector index can reduce available cache space, ultimately degrading LLM throughput, as illustrated in Figure 4 (right).
Beyond memory capacity, GPU retrieval also incurs overheads in compute resource scheduling. Shared memory is used to stage partial distance lookup tables, and each query–cluster pair typically maps to a thread block. As the number of probed clusters increases, so does the occupancy and scheduling pressure on the GPU, further impacting performance.
@@ III-B Opportunity of Tiered Search Structure
Embedding access patterns in recommendation systems are also known to exhibit significant skew, where a small subset of items or users dominates embedding lookup frequency. This observation has motivated several tiered architecture designs that prioritize popular embeddings for faster access  [1, 2, 18, 27].
Inspired by this insight, our work applies a similar idea of tiered acceleration to vector similarity search. However, a key distinction lies in the granularity of retrieval: while recommendation systems typically involve individual embedding lookups, vector search systems such as those based on IVF retrieve and compare distances across hundreds to thousands of vectors per query. Thus, skew in our setting emerges more prominently at the cluster level rather than the vector level.
Consequently, although both domains benefit from tiered designs, the unit of optimization and the manifestation of skew differ substantially. Our approach focuses on cluster-level skew in large-scale retrieval workloads, complementing prior work in recommendation systems.
Figure 6: Hit rates vary significantly across queries, and even when the average hit rate is high, long-tail queries with low hit rates persist. This is particularly evident in the ORCAS dataset.
@@ III-C Variance of Hit Rate across Queries
While tiered resource allocation strategies can accelerate vector search by caching frequently accessed clusters, their effectiveness in deployment is often hindered by query-level variance in hit rates. In particular, long-tail queries with low cache coverage can significantly limit the overall performance gains.
@@
This variance introduces a deployment challenge. Since vector search throughput scales with batch size, retrievers are typically deployed with batching enabled. However, in the presence of low-hit queries within a batch, the entire batch’s processing time is effectively bounded by the slowest query. As a result, even if the average per-query latency is reduced by GPU acceleration, end-to-end latency improvements are constrained.
To fully realize the benefits of tiered or cached retrieval in real-world deployments, it is essential to account for such hit rate variance and long-tail behavior during system design.
@@
Figure 7: System architecture of VectorLiteRAG. The system consists of two stages: offline Hybrid Index Construction (left) and runtime Distributed VectorLiteRAGPipeline (right). During index construction, profiling is conducted to model search latency, query access patterns, and LLM throughput. These are used in the latency-bounded partitioning module to determine the optimal cache size and index partitioning point. The index splitter then constructs sharded GPU/CPU indices and mapping tables. At runtime, queries are processed by a CPU-based coarse quantizer and routed to appropriate workers using mapping tables. Hot clusters are scanned on GPUs, while cold clusters are processed by CPU SIMD cores. A dynamic dispatcher collects early-finished queries and forwards them to LLM workers to reduce tail latency and improve batching efficiency.
In summary, while GPU-based retrieval can vastly outperform CPU methods, it introduces a resource contention problem when co-located with LLMs, due to limited GPU memory and compute capacity. Meanwhile, query access patterns exhibit strong skew: a small fraction of clusters account for the majority of retrieval traffic, making selective caching and tiered search strategies highly effective. However, significant variance in hit rates across queries especially the presence of long-tail queries, poses a major challenge in latency-sensitive deployments, as batching magnifies the performance bottleneck introduced by slow queries. These insights motivate VectorLiteRAG, a system that adaptively partitions the index across GPU and CPU tiers, accounting for workload skew, hit rate variance, and end-to-end latency constraints to optimize both throughput and responsiveness.
@@ IV  VectorLiteRAG
VectorLiteRAG is an optimized RAG system that determines the optimal configuration for a CPU-GPU hybrid vector index. Given the latency constraint, LLM specification, vector index, and system configuration, it computes a partitioning point for tiered search, constructs the corresponding hybrid index, and serves inference requests on a tailored execution pipeline. Figure 7 illustrates the system architecture. VectorLiteRAG consists of two main stages: offline hybrid index construction and the runtime distributed pipeline.
@@
Offline Hybrid Index Construction. This stage begins by profiling CPU-based vector search latency and the query-to-cluster access patterns. It also measures the standalone throughput capacity of the LLM to model contention dynamics. These profiles are used to build a performance model and a cache coverage estimator. Then, based on this model, a latency-bounded partitioning algorithm determines the optimal coverage of ”hot” clusters to be placed on the GPU. The index is then partitioned accordingly, creating sharded sub-indexes for the CPU and GPU.
Distributed VectorLiteRAG Pipeline. At runtime, batched retrieval requests are routed to the appropriate CPU or GPU shards using mapping tables created during index splitting. This reduces the memory and compute pressure on each worker shard and mitigates resource contention with co-located LLM inference. To improve batching efficiency and reduce tail latency, we employ a dynamic dispatcher that proactively identifies early completed queries and forwards them to the next stage of the RAG pipeline.
V Hybrid Index Construction
V-A Profiling-based Performance Modeling

### v3

@@ Abstract
[…]
@@ I Introduction
Retrieval-Augmented Generation (RAG) is a powerful system in natural language processing, particularly for domain-specific question answering and information retrieval tasks [22, 10, 33, 8, 7]. Its key strength lies in combining parametric memory, encoded in the weights of a large language model, with non-parametric memory retrieved from an external knowledge corpus. Although parametric memory provides strong generalization, it is expensive to train and difficult to update. To mitigate this, RAG pipelines first perform similarity search using approximate nearest neighbor search (ANNS) algorithms to retrieve relevant documents from a large database. The retrieved documents are then fed into the LLM’s context to generate up-to-date and reliable responses.
@@
RAG frameworks [20, 24, 38] typically adopt heterogeneous hardware configurations, where vector retrieval is executed on CPUs and LLM generation is served by GPUs. This is driven by the system characteristics: LLM inference requires massive matrix multiplications and benefits significantly from GPU acceleration, whereas retrieval has traditionally been seen as a lighter task suited for CPUs. Offloading retrieval to CPUs allows GPUs to be dedicated to the more compute-intensive generation phase. CPU-based vector search may be sufficient for small vector databases, however, as the dimensionality of the embeddings and the size of the dataset grow, retrieval becomes increasingly compute- and memory- bound. CPUs, with limited parallelism, narrower vector units, and lower memory bandwidth, struggle to handle high-throughput similarity search at scale.
@@
Although the retrieval operation is computationally lighter than the generation phase, it can still benefit significantly from GPU acceleration for two reasons: (1) GPUs feature wide and powerful vector units that enable highly parallelized distance computations, offering superior performance for similarity calculations on long embedding vectors. (2) The retrieval process involves scanning intermediate distance tensors to identify the closest data points in the vector space. These operations are typically implemented as memory lookups a task where GPUs outperform CPUs due to their vectorized memory access and higher I/O.
@@
These compute and memory pressures create a resource tension between the retrieval and generation stages, especially as the vector database grows and CPU-based search fails to meet strict latency requirements. GPU memory is already constrained, with most of it reserved for model weights and KV cache for the LLM. Naively sharding the vector index across all GPUs can lead to memory contention and reduced overall throughput. Alternatively, assigning a disaggregated GPU for retrieval can prevent direct interference between stages, but degrades overall system throughput by reducing the number of available LLM instances, in particular when models require multiple GPUs, enforcing rigid allocation schemes.
Motivated by these challenges, this work explores a holistic approach to optimizing distributed RAG pipelines through joint resource allocation between vector search and LLM generation. We present VectorLiteRAG, a system that partitions the vector index between GPU and CPU-based on query access patterns and LLM deployment configurations, aiming to maximize throughput while meeting latency targets by exploiting the compute power of GPUs across both stages of the RAG pipeline. By analytically modeling similarity search latency, we determine the smallest index portion that needs to be placed on the GPU to satisfy the latency requirement under a given system configuration. Accordingly, VectorLiteRAG offers a latency-aware, throughput-optimized solution that requires no additional hardware resources. This approach is grounded in two key insights:
Access-Skew-Aware Data Layout. VectorLiteRAG leverages a key characteristic of Inverted File (IVF) based retrieval systems [46], that query accesses exhibit skew across clusters. To take advantage of this, VectorLiteRAG incorporates an analytical model that determines the optimal partitioning point and corresponding layout for a multi-GPU system. While the coarse quantizer and cold clusters remain on the CPU, a small subset of hot clusters are cached and distributed across GPUs. The system allocates just enough hot clusters to the GPUs, avoiding both oversubscription of GPU resources during retrieval.
Inter/Intra-Query Variance-Aware Routing. When hot clusters are distributed across GPUs, hit rates vary both across queries (inter-query variance) and across device shards within a query (intra-query variance). Existing systems that enforce fixed retrieval configurations across devices fail to account for this variability and often over-allocate GPU threads. VectorLiteRAG introduces query- and shard- aware routing to avoid such inefficiencies. After determining the most relevant clusters, it dispatches work to CPU or GPU based on their actual expected contribution. It also monitors per-query progress, forwarding early-finishing queries to reduce straggler-induced delays and improve batching efficiency.
@@
  • Access-skew modeling and hit-rate estimation. We characterize access skew in IVF-based retrieval systems and develop a hit-rate estimation method based on observed cluster access patterns.
  • Analytical latency model and SLO-aware partitioning. We construct a latency model that accounts for inter-query variance and use it to determine the optimal CPU-GPU index partitioning point that meets latency targets.
  • Distributed retrieval pipeline. We design a distributed retrieval pipeline that adaptively allocates search tasks across CPUs and GPUs by exploiting inter-device hit rate variance, improving efficiency and avoiding unnecessary GPU resource usage.
@@ II Retrieval Augmented Generation
In a RAG system, user queries are first transformed into vector embeddings using embedding models [34, 30, 42, 36]. These embeddings capture the semantics of the input and enable similarity search by comparing query vectors to a vector database constructed from the knowledge corpus, typically encoded using the same embedding model. State-of-the-art embedding models produce vectors of several thousand dimensions for higher quality, but this increased dimensionality raises the cost of distance computations.
Since exhaustive pairwise search is computationally infeasible at scale, large vector retrieval relies on approximate nearest neighbor search to efficiently identify relevant documents. The retrieved vectors are mapped back to their corresponding documents, which are provided as additional context to the LLM alongside the original query.
@@ II-A Inverted List Index IVF
In contrast, the Inverted File (IVF) index [46] organizes the index as a hierarchical clustering structure. A subset of vectors is first clustered via K-Means to obtain centroids. Then, each database vector is assigned to the closest centroid, forming an inverted list. This structure narrows the search space using only centroid metadata, resulting in low memory overhead and high scalability. As such, IVF is widely adopted and studied in retrieval systems for large knowledge corpora [11, 35, 43, 6, 21, 14]. To further reduce memory usage, quantization techniques are applied on top of IVF. Scalar quantization (SQ) reduces each vector element to a smaller numerical type (e.g., float32 to int8), offering simplicity but limited compression. For higher compression ratios, product quantization (PQ) [13] is commonly used.
@@ II-B Search Operation in IVF Index
Next, a distance lookup table is constructed. Since each vector is quantized into discrete sub-vector codes, each code maps to a representative value, trained and stored in the code-book. By pre-computing distances between the query vector and these representative values, the system avoids computing full distances to every vector. During the scan stage, these LUTs are used to accumulate approximate distances and retrieve the top-k nearest vectors.
@@
Motivated by their superior latency-performance trade-off, we adopt fast scanning in our system to enable efficient and low-latency vector retrieval. However, despite the SIMD capabilities of modern CPUs, CPU-based search can still become a bottleneck, ultimately degrading the responsiveness of the end-to-end RAG system.
@@ III Challenges and Opportunities in RAG serving
While fast scan indexes significantly improve the latency of vector similarity search on CPUs, GPU-based retrieval can offer even greater speedups, due to their wider vector processing units and higher memory bandwidth. As shown in Figure 4 (left), GPU-accelerated IVF search can outperform fast scan methods by nearly an order of magnitude.
Thus, offloading retrieval to the GPU can offer higher speedups for large-scale vector databases where CPU-based search remains a bottleneck. However, this comes with a fundamental trade-off: GPU memory is already heavily utilized by LLMs, particularly for storing KV cache and model weights. Allocating additional memory for the vector index can reduce available cache space, ultimately degrading LLM throughput, as illustrated in Figure 4 (right).
Beyond memory capacity, GPU retrieval additionally incurs scheduling overheads due to increased contention for compute resources. Shared memory is used to stage partial distance lookup tables, and each query–cluster pair typically maps to a thread block. As the number of probed clusters increases, so does the occupancy and scheduling pressure on the GPU, further impacting performance.
@@ III-B Opportunity of Tiered Search Structure
Embedding access patterns in recommendation systems are also known to exhibit significant skew, where a small subset of items or users dominates embedding lookup frequency. This observation has motivated several tiered architecture designs that prioritize popular embeddings for faster access [1, 3, 19, 29, 25, 2, 25]. Inspired by this insight, our work offers tiered acceleration to vector similarity search. However, a key distinction lies in the granularity of memory accesses. In recommendation systems, embedding look-ups are performed via embedding IDs. In contrast, vector similarity search systems conduct fully content-based retrieval, where relevant vectors must be located by computing distances to hundreds or thousands of candidates per query. To identify the nearest vector, the search must access not only the target vector but also neighboring vectors within the cluster.
Moreover, even if each embedding is uniformly accessed, clusters can contain varying numbers of vectors, exacerbating the access skew. This imbalance causes certain clusters to dominate query traffic, creating hot regions in memory access. As a result, skew in our setting emerges more prominently at the cluster level rather than the vector level.
Consequently, although both domains benefit from tiered designs, the unit of optimization and the manifestation of skew differ substantially. Our approach explicitly targets cluster-level skew in large-scale retrieval workloads, enabling effective tiered placement and latency-aware resource allocation that are not directly addressed by prior embedding-centric designs.
Figure 6: Violin plot of hit rate distribution at different cache-coverages. The width of the violin indicates the density of queries with similar hit rates, while the white dot and black bar denote the median and inter-quartile range, respectively. This highlights that increasing cache coverage improves overall hit rates but does not eliminate tail queries with poor hit rates.
@@ III-C Variance of Hit Rate across Queries
While tiered resource allocation strategies can accelerate vector search by caching frequently accessed clusters, their effectiveness in deployment is often hindered by query-level variance in hit rates. Long-tail queries with less cache hits can significantly limit the overall performance gains.
@@
This variance introduces a deployment challenge. Since vector search throughput scales with batch size, retrievers are typically deployed with batching enabled. However, in the presence of low-hit queries within a batch, the entire batch’s processing time is effectively bounded by the slowest query. As a result, even if the average per-query latency is reduced by GPU acceleration, end-to-end latency improvements are constrained. Therefore, to fully realize the benefits of tiered or cached retrieval in real-world deployments, it is essential to account for such hit rate variance and long-tail behavior during system design.
@@
Figure 7: System architecture of VectorLiteRAG. The system has two stages, Left: offline hybrid index construction and Right: runtime distributed pipeline. Profiling guides latency-bounded partitioning to determine cache size and split point, producing sharded indices and mapping tables. At runtime, queries are routed via coarse quantizer and mapping tables, hot clusters run on GPUs, cold clusters on CPUs. A dynamic dispatcher forwards early-finished queries to LLM workers in a timely manner. Blue trails and boxes indicate runtime index refresh and update procedures.
In summary, while GPU-based retrieval can vastly outperform CPU methods, it introduces a resource contention problem when co-located with LLMs, due to limited GPU memory and compute capacity. Meanwhile, query access patterns exhibit strong skew: a small fraction of clusters account for most retrieval traffic, making selective caching and tiered search strategies effective. However, significant variance in hit rates across queries, especially long-tail queries, poses a major challenge in latency-sensitive deployments, as batching magnifies the bottleneck introduced by slow queries. These insights motivate the design of VectorLiteRAG, which adaptively partitions the index across GPU and CPU tiers, accounting for workload skew, hit rate variance, and end-to-end latency constraints to optimize throughput and responsiveness.
@@ IV  VectorLiteRAG
VectorLiteRAG is an optimized RAG system that determines the optimal configuration for a CPU–GPU hybrid vector index. It is organized around tightly integrated components: (1) performance modeling and latency-aware hybrid index construction, and (2) a distributed runtime pipeline for inference serving. Given the latency constraint, LLM, index, and system configuration, VectorLiteRAG computes a partitioning point for tiered search, constructs the hybrid index, and serves inference requests through a tailored pipeline.
@@
Hybrid Index Construction. The first component of VectorLiteRAG focuses on understanding the performance characteristics of the underlying system. This stage profiles CPU-based search latency, query-to-cluster access patterns, and standalone LLM throughput to characterize contention between retrieval and generation. These measurements drive a performance model and cache-coverage estimator, enabling a latency-bounded partitioning algorithm to select hot clusters. The hot clusters are then sharded into GPU sub-indexes.
Distributed VectorLiteRAG Pipeline. The second component is the runtime pipeline that operationalizes the hybrid index. At runtime, batched queries are routed to CPU or GPU shards using mapping tables generated during index construction, allowing each shard to operate with a flexible nprobe budget and reducing contention with LLM. A dynamic dispatcher further improves batching efficiency by advancing early-completing queries to mitigate tail latency.
The partitioning scheme and runtime pipeline are independent of the distance metric or compression method. As long as the index exhibits clustered structure and benefits from GPU acceleration, VectorLiteRAG can identify an effective hybrid configuration and deliver SLO-compliant RAG service.
IV-A Hybrid Index Construction
IV-A1 Profiling-based Performance Modeling